OpenStudy (anonymous):

Multiply and simplify: radical 3(4+radical 12)

OpenStudy (anonymous):

\[\sqrt{12}=\sqrt{4\times 3}=\sqrt{4}\times \sqrt{3}=2\sqrt{3}\]

OpenStudy (anonymous):

so start with \[3(4+2\sqrt{3})\] and use the distributive law . ok?

OpenStudy (anonymous):

okay, so it would be 12+6 radical 3 ?

OpenStudy (anonymous):